The sum of three consecutive odd numbers, the smaller of which is equal to 2n - 1.
Write as an equation

Answer by ikleyn(52800)

.
S = (2n-1) + (2n+1) + (2n+3) = 6n+3 = 3*(2n+1).


So, this sum is three times the middle of the three numbers.